Landon Wallin is an auto mechanic who wishes to start his own business. He will need​ $3800 to purchase tools and equipment. Landon decides to finance the purchase with a 48​-month fixed installment loan with an APR of 5.5​%.

​a) Determine​ Landon's finance charge.

​b) Determine​ Landon's monthly payment.
Answer #1
a) Finance charge $     441.95
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.004583)^-48)/0.004583 i 5.5%/12 = 0.004583
= 42.999114 n 48
Monthly payment = Amount borrowed   / Present value of annuity of 1
= $ 3,800.00 / 42.999114
= $       88.37
Finance charge = Amount repaid - Amount borrowed
= (48*$88.37) - $ 3,800.00
= $     441.95
b) Monthly payment $       88.37